OpenStudy (erinkb99):

Find F ′(x) for F(x)=[2, x^3] sin(t^4) dt

OpenStudy (erinkb99):

\[\int\limits_{x^3}^{2} \sin (t^4) dt\]

jimthompson5910 (jim_thompson5910):

You'd use the fundamental theorem of calculus. When you differentiate an integral, it effectively goes away. That's because integrals are the inverse of a derivative. You'll be left with sin(t^4). Replace each 't' with x^3. Then simplify. Don't forget to multiply by the derivative of x^3 (basically using the chain rule)

OpenStudy (erinkb99):

Possible answers: sin(2^4) – sin(x^12) sin(x^7) –sin(x^12) –3x^2sin(x^12)

jimthompson5910 (jim_thompson5910):

And the result will be negative since the x^3 portion is in the lower limit

OpenStudy (erinkb99):

Ok so sin(t^4) sin((x^3)^4) = sin(x^12)

jimthompson5910 (jim_thompson5910):

that's part of it

OpenStudy (erinkb99):

then sin(x^12)*(3x)?

jimthompson5910 (jim_thompson5910):

3x^2

OpenStudy (erinkb99):

yes, oops

jimthompson5910 (jim_thompson5910):

–3x^2sin(x^12) would be the answer if the x^3 was the upper limit, then the result wouldn't have a negative out front

OpenStudy (erinkb99):

Ok thanks so much!

jimthompson5910 (jim_thompson5910):

sure thing

Nnesha (nnesha):

so first replace t with x^3 \[\sin((x^3)^4)\] and then apply the chain rule ? so why it is not \[\sin ( x^{12}) \cdot 12x^{11}\] ??

jimthompson5910 (jim_thompson5910):

because we're applying the derivative to x^3

Nnesha (nnesha):

ohh okay so multiply the function by the derivative of x^3

Nnesha (nnesha):

ahh okay got it not even applying the chain rule thanks!